OpenStudy (anonymous):

Ralph wants to find the solution to a system of equations using y = x - 4 and y = x2 + 2x - 46. Molly says that Ralph can solve x2 + x - 42 = 0 to find the x coordinates of the solutions to his system. Explain and demonstrate why Molly is correct. @halorazer @PhysicsGuru

OpenStudy (anonymous):

Not going to give you the answer, we you have two equation you can add,subtract,multiply them .... try one of these operations and you will find the answer

OpenStudy (anonymous):

-7?

OpenStudy (anonymous):

You have to perform one of these operations to prove Molly is correct so you are looking to get x2 + x -42 = 0

OpenStudy (anonymous):

What he's saying is what do you do with x-4 and x^2+2x-46 to reach x^2+x-42.

OpenStudy (anonymous):

so im going to use subsituition is that alright

OpenStudy (anonymous):

i dont get it what do you have to do to get that?

OpenStudy (anonymous):

Try what you just stated

OpenStudy (anonymous):

now im confused there no Y in the equation

OpenStudy (anonymous):

that is fine you are proving this equation x2 + x -42 = 0

OpenStudy (anonymous):

yeah i know but how would i write it

OpenStudy (anonymous):

Look at how it is written and what you have now after the substitution.. you want to setup both sides to look like the above equation

OpenStudy (anonymous):

ohhhhhhhhhhhhh so i have to elminate the Y variable?!!?!?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x - 4=x2 + 2x - 46 is this how it would look

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what do you call this method?

OpenStudy (anonymous):

@PhysicsGuru

OpenStudy (anonymous):

You substituted for y with x-4

OpenStudy (anonymous):

so do i move everything to the left or right?

OpenStudy (anonymous):

You have to try what you think and stop looking at others for step by step for all problems because the test later will hurt your grade then.... try it!!

OpenStudy (anonymous):

Guru I know this is why Im practicing with you. im going to move it to the right alright? i jsut need you to tell me if thats right so i dont wast time thats all :)

OpenStudy (anonymous):

It is not a waste of time if you get the wrong answer ... that is learning!! however if you look to others for clarification then you will always doubt yourself. I am a teacher and this is what students will do because they are afraid of incorrect answers.

OpenStudy (anonymous):

Thanks! ok i will let you know how far i get

OpenStudy (anonymous):

x=x^2 + 2^x - 42

OpenStudy (anonymous):

@PhysicsGuru

OpenStudy (anonymous):

do we make those thingys with two equations?

OpenStudy (anonymous):

factor i mean?

OpenStudy (anonymous):

@nincompoop pls help me?

OpenStudy (nincompoop):

sure

OpenStudy (nincompoop):

do you want from the very basic idea of systems of linear equations?

OpenStudy (anonymous):

yeah

OpenStudy (nincompoop):

when two lines are involved in a cartesian graph, we can make certain predictions at what point those two lines may intersect or not intersect at all.

OpenStudy (nincompoop):



OpenStudy (nincompoop):